# On decoding extended Han-Zhang codes

### v1

@@ II-C ECPs and deep holes
( [10, 36]) Let $\mathcal{C}$} be an $n,k]_{q}$ linear code. Let $\mathcal{A}$ and $\mathcal{B}$ be two linear codes with the same length $n$ over $\mathbb{F}_{q}$. We call $(\mathcal{A},\mathcal{B})$ an $\ell$-error-locating pair ($\ell$-ELP) of $\mathcal{C}$} if all of following three conditions hold:
${\rm i)}~\mathcal{A}\star\mathcal{B}\subseteq{\mathcal{C}}^{\perp};~~{\rm ii)}~d(\mathcal{B}^{\perp})>\ell;~~{\rm iii)}~\dim(\mathcal{A})>\ell.$
Moreover, if an $\ell$-ELP $(\mathcal{A},\mathcal{B})$ satisfies the additional condition
${\rm iv)}~d(\mathcal{A})+d({\mathcal{C}})>n,$
we call it an $\ell$-error-correcting pair ($\ell$-ECP) of $\mathcal{C}$}.
@@
( [17, Remark 2.1]) Let ${\mathcal{C}}_{i}$ be an $n,k_{i}]_{q}$ linear code for $i=1,2$ with ${\mathcal{C}}_{1}\subseteq{\mathcal{C}}_{2}$ and $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of ${\mathcal{C}}_{2}$. Then the following statements hold.
@@
( [34, Example 4.2] and [34, Theorem 6.2]) Let $\mathcal{C}$} be an $n,n-2\ell,2\ell+1]_{q}$ MDS code, where $\ell$ is a positive integer with $1\leq\ell\leq\frac{n}{2}-1$. Then $\mathcal{C}$} has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ if and only if $\mathcal{C}$} is a GRS code. Moreover, both $\mathcal{A}$ and $\mathcal{B}$ are GRS codes with the same evaluation-point sequence as that of $\mathcal{C}$}, and $\mathcal{B}=(\mathcal{A}\star\mathcal{C})^{\perp}$.
@@
( [37, Proposition 2.5]) Let $\mathcal{C}$} be an $n,n-2\ell,2\ell+1]_{q}$ MDS code and let $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of $\mathcal{C}$}. Then $\mathcal{A}$ is an $n,\ell+1,n-\ell]_{q}$ MDS code.
Definition 6.
( [21]) Let $\mathcal{C}$} be an $n,k]_{q}$ linear code. For any vector $\mathbf{u}\in\mathbb{F}_{q}^{n}$, its error distance to $\mathcal{C}$} is defined as $d(\mathbf{u},{\mathcal{C}})=\min\{d(\mathbf{u},\mathbf{c}):~\mathbf{c}\in{\mathcal{C}}\}$. We call $\rho({\mathcal{C}})=\max\{d(\mathbf{u},{\mathcal{C}}):~\mathbf{u}\in\mathbb{F}_{q}^{n}\}$ the covering radius of $\mathcal{C}$} and the deep holes of $\mathcal{C}$} are just the vectors in $\mathbb{F}_{q}^{n}$ that achieve $\rho({\mathcal{C}})$.
The following lemma provides a way to get more non-GRS MDS codes from a known non-GRS MDS code with respect to some conditions.
Lemma 9.
( [33, Lemma 2.4]) Let $\mathcal{C}$} be an $n+1,k,n-k+2]_{q}$ MDS code with generator matrix $G$ and covering radius $\rho({\mathcal{C}})=n-k+1$. Then for any $\mathbf{u}\in\mathbb{F}_{q}^{n+1}$, $\mathbf{u}$ is a deep hole of $\mathcal{C}$} if and only if the matrix $\left(\left(\begin{array}[]{c}G\\ \mathbf{u}\end{array}\right),~\infty_{k+1}\right)$ generates an $n+2,k+1,n-k+2]_{q}$ MDS code $\mathcal{C}$$^{\prime}$. Moreover, if $\mathcal{C}$} is non-GRS, so is $\mathcal{C}$$^{\prime}$.
The following two lemmas are also relevant in identifying deep holes of extended Han-Zhang codes.
Lemma 10.
([1, Lemma 7.3]) Let $\mathcal{C}$} be an $n,k]_{q}$ linear code with generator matrix $G$. Then $\mathcal{C}$} is MDS if and only if any $k$ columns of $G$ are linearly independent.
Lemma 11.
( [51, Lemma 2.3]) Let $m$ be a fixed positive integer and let $I_{m}=\left\{0,\ldots,m-1\right\}=\left\{t_{1},\ldots,t_{s}\right\}\cup\left\{r_{1},\ldots,r_{s^{\prime}}\right\}$ be any partition of $I_{m}$ with $m=s+s^{\prime}$, $0=t_{1}<\ldots<t_{s}=m-1$ and $r_{1}<\ldots<r_{s^{\prime}}$. Then for any ${\mathcal{S}}=\left\{a_{1},\ldots,a_{s}\right\}\subseteq\mathbb{F}_{q}$, we have
$\det\left(\begin{array}[]{ccc}a_{1}^{t_{1}}&\ldots&a_{s}^{t_{1}}\\ a_{1}^{t_{2}}&\ldots&a_{s}^{t_{2}}\\ \vdots&\ddots&\vdots\\ a^{t_{s}}_{1}&\ldots&a^{t_{s}}_{s}\\ \end{array}\right)=\prod_{1\leq i<j\leq s}(a_{j}-a_{i})\cdot\det\left(\begin{array}[]{ccc}\sigma_{s-r_{1}}({\mathcal{S}})&\ldots&\sigma_{s-r_{s^{\prime}}}({\mathcal{S}})\\ \sigma_{s-r_{1}+1}({\mathcal{S}})&\ldots&\sigma_{s-r_{s^{\prime}}+1}({\mathcal{S}})\\ \vdots&\ddots&\vdots\\ \sigma_{s-r_{1}+s^{\prime}-1}({\mathcal{S}})&\ldots&\sigma_{s-r_{s^{\prime}}+s^{\prime}-1}({\mathcal{S}})\\ \end{array}\right).$
III ECPs of extended Han-Zhang codes
From now on, we explore the answers to Problem 1. In this section, we study the existence and specific forms of ECPs of extended Han-Zhang codes, and further provide a decoding algorithm for extended Han-Zhang codes based on their ECPs.
III-A The existence of $\ell$-ECPs of extended Han-Zhang codes
First of all, we note that the following nested relationships between extended Han-Zhang codes and (extended) GRS codes.
@@ III-A The existence of $\ell$-ECPs of extended Han-Zhang codes
Additionally, for any $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$ (such $\gamma$ exists since $n\leq q-1$), it follows from [37] and [56, Lemma 4] that there exists a set ${\mathcal{S}}^{\prime}=\{a_{1}^{\prime},\ldots,a_{n+1}^{\prime}\}\subseteq\mathbb{F}_{q}$ and a vector ${{\mathbf{v}}}^{\prime}=(v_{1}^{\prime},\ldots,v_{n+1}^{\prime})\in(\mathbb{F}_{q}^{*})^{n+1}$ such that
@@
with
$a_{i}^{\prime}=\left\{\begin{array}[]{ll}(a_{i}-\gamma)^{-1},&{\rm if}~1\leq i\leq n,\\ 0,&{\rm if}~i=n+1,\end{array}\right.~{\rm and}~v_{i}^{\prime}=\left\{\begin{array}[]{ll}v_{i}(a_{i}-\gamma)^{k},&{\rm if}~1\leq i\leq n,\\ 1,&{\rm if}~i=n+1.\end{array}\right$.
@@
In the following, we investigate the existence of $\ell$-ECPs of extended Han-Zhang codes and give explicit forms of these $\ell$-ECPs if they exist.
@@
    If $2\nmid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ does not have an $\ell$-ECP. Moreover, ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $(\ell-1)$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf 1})~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{{\mathbf{u}}}^{\prime}),$
@@
    If $2\mid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1).$
@@
1) Since $\mathcal{S}$} is $k$-zero-sum free, it follows from Lemma 4.1) that $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+2$. From $2\nmid(n-k)$ and $3\leq k\leq n-2$, we get that $3\leq k\leq n-3$ and $2\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k+1}{2}\leq\frac{n-2}{2}<\frac{n+1}{2}-1$. Combining Lemmas 4.1) and 5, we immediately get that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)={\mathcal{C}}_{n-2\ell+1}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,n-2\ell+1,2\ell+1]_{q}$ non-GRS MDS code. By Lemma 7, if ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP, then it is a GRS code, a contradiction. Hence, ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ does not have an $\ell$-ECP.
Moreover, when ${\mathcal{S}}=\mathbb{F}_{q}$, it follows from [13, Corollary 1] that $\mathcal{S}$} is $k$-zero-sum free if and only if $q$ is even and k=q2k=q-2, which contradicts the fact $3\leq k\leq n-3$. Therefore, we conclude that $n\leq q-1$ and $\mathbb{F}_{q}\setminus{\mathcal{S}}\neq\emptyset$ under the given conditions. It then follows from Lemma 12 that
${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)={\mathcal{C}}_{n-2\ell+1}({\mathcal{S}},{\bf v},\infty)\subseteq{\mathrm{GRS}}_{n-2\ell+2}({\mathcal{S}},{\bf v},\infty)\subseteq{\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime}),$ (14)
where ${\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime})$ is an $n+1,n-2\ell+3,2\ell-1]_{q}$ GRS code with ${\mathcal{S}}^{\prime}=\{(a_{1}-\gamma)^{-1},\ldots,(a_{n}-\gamma)^{-1},0\}\subseteq\mathbb{F}_{q}$, ${{\mathbf{v}}}^{\prime}=(v_{1}(a_{1}-\gamma)^{k},\ldots,v_{n}(a_{n}-\gamma)^{k},1)\in(\mathbb{F}_{q}^{*})^{n+1}$, and any $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$. Note that $2\ell-1=2(\ell-1)+1$ and $1\leq\ell-1<\frac{n+1}{2}-1$, then it follows from Lemmas 1, 7, and 8 that ${\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime})$ has an $(\ell-1)$-ECP $(\mathcal{A}^{\prime},\mathcal{B}^{\prime})$, where $\mathcal{A}^{\prime}={\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf r})$ is an $n+1,\ell,n-\ell+2]_{q}$ MDS code for some ${\bf r}\in(\mathbb{F}_{q}^{*})^{n+1}$ and
$\mathcal{B}^{\prime}=({\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf r})\star{\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime}))^{\perp}={\mathrm{GRS}}_{n-\ell+2}({\mathcal{S}}^{\prime},{\bf r}\star{\bf v}^{\prime})^{\perp}={\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf r}^{-1}\star{\bf v}^{-1}\star{{\mathbf{u}}}^{\prime})$
@@
$=$ $\left\{\begin{array}[]{ll}(a_{i}-\gamma)^{n}\prod_{1\leq j\neq i\leq n}(a_{j}-\gamma)(a_{j}-a_{i})^{-1},&{\rm if}~1\leq i\leq n,\\ (-1)^{n}\prod_{1\leq j\leq n}(a_{j}-\gamma),&{\rm if}~i=n+1.\end{array}\right$.
@@
$(\mathcal{A},\mathcal{B})=\left({\bf r}^{-1}\star\mathcal{A}^{\prime},{\bf r}\star\mathcal{B}^{\prime}\right)=\left({\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf 1}),{\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{{\mathbf{u}}}^{\prime})\right)$
is an $(\ell-1)$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof of 1).
2) In this case, we have that $1\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k}{2}\leq\frac{n-3}{2}<\frac{n+1}{2}-1$ and ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has parameters $n+1,n-2\ell,2\ell+2]_{q}$. Since a parity-check matrix of a linear code is a generator matrix of its dual code, it then follows from Lemma 3 that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$ has a generator matrix of the form
@@
$G_{\mathcal{A}}=\begin{pmatrix}1&\ldots&1&0\\ a_{1}&\ldots&a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ a^{\ell-1}_{1}&\ldots&a^{\ell-1}_{n}&0\\ a^{\ell}_{1}&\ldots&a^{\ell}_{n}&1\end{pmatrix}~{\rm and}~G_{\mathcal{B}}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-1}&-1\\ \end{pmatrix}$,
@@
$\begin{split}d(\mathcal{B}^{\perp})&=d({\mathrm{GRS}}_{n+1-\ell}({\mathcal{S}},{\bf v},\infty))=n+1-(n+1-\ell)+1=\ell+1>\ell,\\ \dim(\mathcal{A})&=\ell+1>\ell,~{\rm and}\\ d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))&=n+1-(\ell+1)+1+2\ell+2=n+\ell+3>n+1,\end{split}$
@@
$(\mathcal{A},\mathcal{B})=\left({\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof of 2). ∎

### v2

@@ II-C ECPs of linear codes
( [8, 29]) Let $\mathcal{C}$} be an $n,k]_{q}$ linear code. Let $\mathcal{A}$ and $\mathcal{B}$ be two linear codes with the same length $n$ over $\mathbb{F}_{q^{t}}$ for a given $t\in\mathbb{N}$. We call $(\mathcal{A},\mathcal{B})$ an $\ell$-error-correcting pair ($\ell$-ECP) of $\mathcal{C}$} over $\mathbb{F}_{q^{t}}$ if all the following conditions hold:
  1. 1.
    $\mathcal{A}\star\mathcal{B}\subseteq{\mathcal{C}}^{\perp}$, with $\mathcal{C}$$^{\perp}$ seen as a linear code over $\mathbb{F}_{q^{t}}$.
  2. 2.
    $d(\mathcal{B}^{\perp})>\ell$.
  3. 3.
    $\dim_{\mathbb{F}_{q^{t}}}(\mathcal{A})>\ell$.
  4. 4.
    $d(\mathcal{A})+d({\mathcal{C}})>n$.
@@
( [14, Remark 2.1]) Let ${\mathcal{C}}_{i}$ be an $n,k_{i}]_{q}$ linear code for $i=1,2$ with ${\mathcal{C}}_{1}\subseteq{\mathcal{C}}_{2}$ and $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of ${\mathcal{C}}_{2}$. Then the following statements hold.
@@
( [28, Example 4.2] and [28, Theorem 6.2]) Let $\mathcal{C}$} be an $n,n-2\ell,2\ell+1]_{q}$ MDS code, where $\ell$ is a positive integer with $1\leq\ell\leq\frac{n}{2}-1$. Then $\mathcal{C}$} has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ over $\mathbb{F}_{q}$ if and only if $\mathcal{C}$} is a GRS code. Moreover, both $\mathcal{A}$ and $\mathcal{B}$ are GRS codes with the same evaluation-point sequence as that of $\mathcal{C}$}, and $\mathcal{B}=(\mathcal{A}\star\mathcal{C})^{\perp}$.
@@
( [30, Proposition 2.5]) Let $\mathcal{C}$} be an $n,n-2\ell,2\ell+1]_{q}$ MDS code and let $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of $\mathcal{C}$}. Then $\mathcal{A}$ is an $n,\ell+1,n-\ell]_{q}$ MDS code.
III The existence and specific forms of ECPs of ESGRS codes
From now on, we explore the answers to Problem 1. In this section, we study the existence and specific forms of ECPs of ESGRS codes in consideration of the MDS and NMDS cases.
III-A The case of MDS ESGRS codes
First of all, we note that the following nested relationships between ESGRS codes and (extended) GRS codes.
@@ III-A The case of MDS ESGRS codes
Additionally, for any $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$ (such $\gamma$ exists since $n\leq q-1$), it follows from [30] that there exists a set ${\mathcal{S}}^{\prime}=\{a_{1}^{\prime},\ldots,a_{n+1}^{\prime}\}\subseteq\mathbb{F}_{q}$ and a vector ${{\mathbf{v}}}^{\prime}=(v_{1}^{\prime},\ldots,v_{n+1}^{\prime})\in(\mathbb{F}_{q}^{*})^{n+1}$ such that
@@
by taking
$a_{i}^{\prime}=\left\{\begin{array}[]{ll}(a_{i}-\gamma)^{-1},&{\rm if}~1\leq i\leq n,\\ 0,&{\rm if}~i=n+1,\end{array}\right$.
and
$v_{i}^{\prime}=\left\{\begin{array}[]{ll}v_{i}(a_{i}-\gamma)^{k},&{\rm if}~1\leq i\leq n,\\ 1,&{\rm if}~i=n+1.\end{array}\right$.
@@
Remark 1.
Note that the parameters and dual properties of SGRS codes, obtained by removing the last coordinate from each codeword of ESGRS codes, have been investigated in [12, 19, 24]. Moreover, SGRS codes have been shown to be either MDS or NMDS codes [12], and in certain cases, they are specifically non-GRS MDS codes [19, 24]. From the proof of Lemma 8, ESGRS codes may coincide with SGRS codes when their lengths are less than $q$ and certain additional conditions are satisfied. Although this possibility cannot be excluded, it is important to note that the decoding problem for SGRS codes also remains open. In fact, if such an equivalence holds, the decoding algorithm we derive in the sequel would be directly applicable to SGRS codes as well.
In the following, we investigate the existence of $\ell$-ECPs of MDS ESGRS codes and give explicit forms of these $\ell$-ECPs if they exist.
@@
    If $2\nmid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ does not have an $\ell$-ECP. Moreover, ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $(\ell-1)$-ECP $(\mathcal{A},\mathcal{B})$ over $\mathbb{F}_{q}$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf 1})$
    and
    $\mathcal{B}={\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{{\mathbf{u}}}^{\prime}),$
@@
    If $2\mid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ over $\mathbb{F}_{q}$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)$
    and
    $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1).$
@@
1) Since $\mathcal{S}$} is $k$-zero-sum free, it follows from Lemma 3.1) that $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+2$. From $2\nmid(n-k)$ and $3\leq k\leq n-2$, we get that $3\leq k\leq n-3$ and $2\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k+1}{2}\leq\frac{n-2}{2}<\frac{n+1}{2}-1$. Combining Lemmas 3.1) and 4, we immediately get that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)={\mathcal{C}}_{n-2\ell+1}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,n-2\ell+1,2\ell+1]_{q}$ non-GRS MDS code. By Lemma 6, if ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP, then it is a GRS code, a contradiction. Hence, ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ does not have an $\ell$-ECP.
Moreover, when ${\mathcal{S}}=\mathbb{F}_{q}$, it follows from [23, Corollary 2.8] that $\mathcal{S}$} always contains a $k$-zero-sum subset if $3\leq k\leq n-3$. Therefore, we conclude that $n\leq q-1$ and $\mathbb{F}_{q}\setminus{\mathcal{S}}\neq\emptyset$, since $\mathcal{S}$} is assumed to be $k$-zero-sum free. It then follows from Lemma 8 that
$\begin{split}{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)&={\mathcal{C}}_{n-2\ell+1}({\mathcal{S}},{\bf v},\infty)\\ &\subseteq{\mathrm{GRS}}_{n-2\ell+2}({\mathcal{S}},{\bf v},\infty)\\ &\subseteq{\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime}),\end{split}$ (7)
where ${\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime})$ is an $n+1,n-2\ell+3,2\ell-1]_{q}$ GRS code with ${\mathcal{S}}^{\prime}=\{(a_{1}-\gamma)^{-1},\ldots,(a_{n}-\gamma)^{-1},0\}\subseteq\mathbb{F}_{q}$, ${{\mathbf{v}}}^{\prime}=(v_{1}(a_{1}-\gamma)^{k},\ldots,v_{n}(a_{n}-\gamma)^{k},1)\in(\mathbb{F}_{q}^{*})^{n+1}$, and any $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$. Note that $2\ell-1=2(\ell-1)+1$ and $1\leq\ell-1<\frac{n+1}{2}-1$, then it follows from Lemmas 1, 6, and 7 that ${\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime})$ has an $(\ell-1)$-ECP $(\mathcal{A}^{\prime},\mathcal{B}^{\prime})$, where $\mathcal{A}^{\prime}={\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf r})$ is an $n+1,\ell,n-\ell+2]_{q}$ MDS code for some ${\bf r}\in(\mathbb{F}_{q}^{*})^{n+1}$ and
$\mathcal{B}^{\prime}$ $=({\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf r})\star{\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime}))^{\perp}$
$={\mathrm{GRS}}_{n-\ell+2}({\mathcal{S}}^{\prime},{\bf r}\star{\bf v}^{\prime})^{\perp}$
$={\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf r}^{-1}\star{\bf v}^{-1}\star{{\mathbf{u}}}^{\prime})$
@@
$=$ $\left\{\begin{array}[]{ll}(a_{i}-\gamma)^{n}\prod_{1\leq j\neq i\leq n}\frac{a_{j}-\gamma}{a_{j}-a_{i}},&{\rm if}~1\leq i\leq n,\\ (-1)^{n}\prod_{1\leq j\leq n}(a_{j}-\gamma),&{\rm if}~i=n+1.\end{array}\right$.
@@
$\left({\bf r}^{-1}\star\mathcal{A}^{\prime},{\bf r}\star\mathcal{B}^{\prime}\right)=\left({\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf 1}),{\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{{\mathbf{u}}}^{\prime})\right)$
is an $(\ell-1)$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ over $\mathbb{F}_{q}$. This completes the proof of 1).
2) In this case, we have that $1\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k}{2}\leq\frac{n-3}{2}<\frac{n+1}{2}-1$ and ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has parameters $n+1,n-2\ell,2\ell+2]_{q}$. Since a parity-check matrix of a linear code is a generator matrix of its dual code, it then follows from (4) that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$ has a generator matrix of the form
@@
$G_{\mathcal{A}}=\begin{pmatrix}1&\ldots&1&0\\ a_{1}&\ldots&a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ a^{\ell-1}_{1}&\ldots&a^{\ell-1}_{n}&0\\ a^{\ell}_{1}&\ldots&a^{\ell}_{n}&1\end{pmatrix}$
and
$G_{\mathcal{B}}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-1}&-1\\ \end{pmatrix}$,
@@
$d(\mathcal{B}^{\perp})=d({\mathrm{GRS}}_{n+1-\ell}({\mathcal{S}},{\bf v},\infty))=\ell+1>\ell,$
$\dim(\mathcal{A})=\ell+1>\ell,$
and
$d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$ $=n+1-(\ell+1)+1+2\ell+2$
$=n+\ell+3$
$>n+1,$
@@
$\left({\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ over $\mathbb{F}_{q}$. This completes the proof of 2). ∎
III-B The case of NMDS ESGRS codes
This subsection focuses on NMDS ESGRS codes. In this case, the existence of $\ell$-ECPs is always guaranteed. The following theorem presents the explicit forms of their $\ell$-ECPs, which are constructed based on suitable variants of GRS codes.